OpenStudy (anonymous):

Choose the correct simplification of the expression (4x^3y^2z^4)(2x^3y^4z^3). 8x^9y^8z^12 6x^6y^6z^7 8x^6y^6z^7 6x^9y^8z^12

OpenStudy (anonymous):

C. 8x^6y^6z^7

OpenStudy (anonymous):

When simplifying expressions with similar variables and powers such as this one, you simply multiply like terms, then add the powers if you are multiplying. In this case, we have the like terms of 4x and 2x. We would simplify this to 8x because 2*4=8, so you just add the x. We have powers of 3y and 3y, so since 3+3=6, your power in your answer would result in 6y. Since 4z+2z=6z because 4+2=6, 6z is your answer for the next exponent, and since to the POWER of 4 plus the power of 3 is the power of 7 because 4+3=7, you would power those numbers and variables by 7.

OpenStudy (anonymous):

Understand it @jaderbrown ?

OpenStudy (anonymous):

yes I do thank you!

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

@AwesomeB copy paste haha

OpenStudy (anonymous):

lol i solved it cause it is easy but since openstudy requires you to show explanation, mine was too long

OpenStudy (uri):

(4x^3y^2z^4)(2x^3y^4z^3). Add the numericals and add all x powers together,y together and z together 5*2 =8, x^3+x^3 =x^6 ,y^2+y^4=y^6,z^4+z^3=z^7 That will be, 8x^6y^6z^7

OpenStudy (anonymous):

uri can yoi help me with my question? i tagged you in it

OpenStudy (uri):

Let me see.